OpenStudy (anonymous):

Can someone please hep me solve Let f(x) = 2x - 6. Solve f-1(x) when x = 2

OpenStudy (zzr0ck3r):

you need to solve this \(2x-6 = 2\)

OpenStudy (zzr0ck3r):

the x in the inverse function is the y in the original function.

OpenStudy (anonymous):

can you show me how

OpenStudy (zzr0ck3r):

add 6 to both sides and what do you have?

OpenStudy (anonymous):

so the answer would be 4

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

can u help me with one more please

OpenStudy (zzr0ck3r):

this question is asking what do you need to put in f(x) to get out 2 i.e. what is f^-1(2)

OpenStudy (zzr0ck3r):

sure

OpenStudy (anonymous):

Let f(x) = x2 - 16. Find f-1(x)

OpenStudy (anonymous):

f-1(x) = 2

OpenStudy (anonymous):

can u show me how you did it please

OpenStudy (zzr0ck3r):

this function does not have an inverse

OpenStudy (anonymous):

?? that was an inverse thought it was a prime

OpenStudy (anonymous):

sorry my bad

OpenStudy (anonymous):

these are the answer choices they gave me 4 square root of x ±square root of x plus 16 x squared over 16 1 over quantity of x squared minus 16

OpenStudy (zzr0ck3r):

\(f(x) = x^2-16\) this function does not have a inverse

OpenStudy (zzr0ck3r):

but the answer they want is the second one

OpenStudy (zzr0ck3r):

but it is wrong lol

OpenStudy (anonymous):

idk the answers i gave you were the answer choices to that equation

OpenStudy (zzr0ck3r):

they are looking for the second answer

OpenStudy (zzr0ck3r):

but this question is flawed and should be removed from the book/source. Tell your teacher that. \(f(x)\) means a function, and thus \(f^{-1}(x)\) must also be a function, but +-sqrt{x+16} is not a function.

OpenStudy (zzr0ck3r):

i.e. f(x) must be onto and one to one, which it is not.

OpenStudy (anonymous):

ok i will and thank you

OpenStudy (zzr0ck3r):

the question should have said, what function models the inverse image of f(x)